OpenStudy (anonymous):

let f(x)=-4x=7 and g(x)= 10x-6. find g(g(x))

OpenStudy (anonymous):

find f(g(x))

OpenStudy (popitree):

f(x) = -4x + 7 f(g(x)) = -4(10x-6)+7 =-40x+31

OpenStudy (anonymous):

Ok so with these composite functions, we simply are going to substitute the equation g(x) into the x of the f(x) function.

OpenStudy (anonymous):

So, go ahead and substitute the function g(x) = 10x - 6 as the "x" in the function f(x).

OpenStudy (anonymous):

Also you have the f(x) function written improperly, the double equals signs are confusing, do you mean f(x) = -4x - 7?

OpenStudy (anonymous):

ya f(x) = -4x - 7 is right, sorry

OpenStudy (anonymous):

Perfect, so show me what f(x) looks like after substituting g(x) into the "x".

OpenStudy (anonymous):

-4(10+6)+7 ....?

OpenStudy (anonymous):

Ok, sorry, one sec.

OpenStudy (anonymous):

Almost right, you forgot the x, so it looks like this, f(g(x)) -> -4(10x - 6) - 7 Now we just evaluate that function by simplifying it.

OpenStudy (anonymous):

thats the part im not getting, the simplify part

OpenStudy (anonymous):

Alright, so I'll walk you through it step by step =0)

OpenStudy (anonymous):